What would the answer be when you evaluate the function sec 11pi/6 without a calculator, and how would you do this?
Answer by lwsshak3(11628)

What would the answer be when you evaluate the function sec 11pi/6 without a calculator, and how would you do this?
**
11π/6 is in quadrant IV where a reference angle of π/6 applies and where sec>0
sec 11pi/6=sec(π/6)=2/√3=(2√3)/3